# Chat2SVG: Vector Graphics Generation with Large Language Models and Image Diffusion Models

### v1

@@ 3.3 Iterative Editing
Our system supports iterative refinement through multiple rounds of natural language instructions. However, across different iterations, when passing the SVG template image through SDEdit, the enhanced images may contain varying decorative details, potentially leading to inconsistencies in the final optimized SVG. To maintain consistency across editing iterations, we instruct the LLM to output both the modified SVG script and a list of specifically changed paths. This precise tracking enables selective optimization of modified shapes while preserving unaltered elements from previous iterations, maintaining visual coherence throughout the entire refinement process.
@@ 4 Experiments
Implementation Details. In our experiments, we use claude-3-5-sonnet as our backbone LLM model due to its leading generation capabilities. To evaluate our approach, we create a prompt set by having the LLM generate text prompts across five categories: animals, food, objects, scenes, and novel concept combinations. For each prompt, we generate one SVG template and perform rounds of visual rectification. This process repeats times, yielding candidate SVG templates per prompt. These templates are rendered as images, and we follow a standard practice to select the highest quality SVG using the ImageReward [45] metric. Detailed optimization configurations are provided in the supplementary materials.
Baselines. We compare our approach against five state-of-the-art image-based text-to-SVG generation methods: CLIPDraw [9], DiffSketcher [43], VectorFusion [16], SVGDreamer [44], and T2V-NPR [50]. Current language-based methods [41, 35] are confined to specific categories, with no support for general text-to-SVG generation, so we do not include them in our comparison. For a fair comparison, we set these baseline methods to use a similar number of shapes as our optimized SVG.
@@
Image-level Fidelity. We evaluate visual fidelity using a ground-truth dataset of well-designed vector graphics, specifically colored SVG files downloaded from SVGRepo111https://www.svgrepo.com. We compute the Fréchet Inception Distance (FID) [13] between the rendered images of our generated SVG and this professional design collection, using features extracted by the CLIP image encoder [28].
Path-level Regularity. We assess path quality using a transformer-based Auto-Encoder trained on the FIGR-8-SVG [6] dataset, following the DeepSVG [4] architecture. Unlike previous methods [50] that need a fixed number of commands, our model encodes each drawing command into a latent vector, accommodating varying command lengths. After training, we compute the mean of command latent embeddings for each path and calculate the FID between these and the ground truth paths from FIGR-8-SVG. This quantifies how well our paths align with professional SVG design patterns.
Text-level Alignment. We assess semantic alignment by computing the CLIP score [28] between the text prompt and the rendered SVG image.
@@ 4.1 Comparison with Existing Methods
| CLIPDraw [9] | |||||||||
| DiffSketcher [43] | |||||||||
| VectorFusion [16] | |||||||||
| SVGDreamer [44] | |||||||||
| T2V-NPR [50] | |||||||||
@@
Qualitative Comparison. In Figure 3, we present a side-by-side comparison between our Chat2SVG and the baselines. Methods that optimize open-ended strokes based on CLIP or Diffusion models (i.e., CLIPDraw [9] and DiffSketcher [43]) produce results with messy visual appearances and poor text alignment. These methods sometimes fail to synthesize complete objects, such as the flamingo (first row) and the plane (third row) in Figure 3. In contrast, our method generates SVG results with clear layouts and greater fidelity to the prompts.
Methods that optimize closed shapes via score distillation of diffusion models (i.e., VectorFusion [16], SVGDreamer [44], and T2V-NPR [50]) can better align with the input text, but still produce fragmented paths with limited semantic meaning. For instance, in the plane example (third row of Figure 3), VectorFusion and SVGDreamer create planes and clouds composed of multiple jagged, irregular, and fragmented shapes, which only appear meaningful when viewed as a whole. This hinders convenient editing by graphic designers. T2V-NPR [50] addresses the issue of jagged paths using a path VAE (although it can be overly smooth at times) and reduces fragmented shapes by merging shapes with similar colors. However, its merging operation ignores the semantic meaning of shapes, resulting in the body of the plane being merged with the surrounding clouds. Additionally, the semantic meaning of the path is sometimes unclear, as exemplified by the cloud in the bottom-right corner, whose contour is actually approximated by the blue background instead of a true cloud shape. In contrast, our Chat2SVG adopts a unique approach by using LLMs to generate paths representing semantic components and performing dual-stage optimization based on image diffusion to enhance path expressiveness, thus ensuring both path regularity and visual fidelity of generated SVG.
@@ 5 Conclusion
Despite our method’s superior results, there are several aspects that could be further improved: (1) LLMs may produce imprecise visual layouts, such as the burger in Figure 7. This can be alleviated by collecting an SVG dataset and extracting the corresponding layouts to fine-tune LLMs. (2) Our approach uses SAM to identify new parts. However, SAM may overlook important regions, like the pink frosting on the donut in Figure 7. Incorporating semantic-aware SAM [19] could potentially produce more accurate results. (3) The pretrained SVG VAE from [50] has a fixed number of commands. This makes it hard to represent intricate shapes. For example, the lion’s mane in Figure 7 lacks sufficient detail. Training a new SVG VAE with varying-length commands would help solve this problem. Our work can be extended to generate other types of vectorized data, such as icons and fonts. Additionally, through consistent editing of SVG across multiple frames, coupled with shape interpolation, our Chat2SVG has the potential to facilitate keyframe animation generation.
@@ Appendix A Overview
More generated and edited SVG examples (Section E) are available on the project website.
@@ Appendix E More Results
To provide a more user-friendly visualization, we have put all generated SVG, including text-guided generation and editing results, on the project page.
| System Prompt |

### v2

@@ 3.3 Iterative Editing
Our system supports iterative refinement through multiple rounds of natural language instructions. However, when passing the edited SVG template image into SDEdit, image diffusion models may introduce different shape variations and decorative details, which compromise the user’s intention to maintain the already optimized shapes. To preserve consistency across editing iterations, we instruct the LLM to output both the modified SVG script and a list of specifically changed paths. This precise tracking enables selective optimization of modified shapes while preserving unaltered elements from previous iterations, maintaining visual coherence throughout the entire refinement process.
@@ 4 Experiments
Implementation Details. In our experiments, we use claude-3-5-sonnet as our backbone LLM model due to its leading generation capabilities. To evaluate our approach, we create a prompt set by having the LLM generate text prompts across five categories: animals, food, objects, scenes, and novel concept combinations. For each prompt, we generate one template and perform rounds of visual rectification. This process is repeated times, yielding candidate SVG templates per prompt. These templates are rendered as images, and we follow a standard practice to select the highest quality SVG using the ImageReward [xu2024imagereward] metric. Detailed optimization configurations are provided in the supplementary materials.
Baselines. We compare our approach against five state-of-the-art image-based text-to-SVG generation methods: CLIPDraw [frans2021clipdraw], DiffSketcher [xing2023diffsketcher], VectorFusion [jain2022vectorfusion], SVGDreamer [xing2024svgdreamer], and T2V-NPR [zhang2024text]. Current language-based methods [wu2023iconshop, tang2024strokenuwa] are confined to specific categories, with no support for general text-to-SVG generation, so we do not include them in our comparison. For a fair comparison, we set these baseline methods to use a similar number of shapes as our optimized SVG.
@@
Image-level Fidelity. We evaluate visual fidelity using a ground-truth dataset of well-designed vector graphics, specifically colored SVG files downloaded from SVGRepo111https://www.svgrepo.com. We compute the Fréchet Inception Distance (FID) [heusel2017gans] between the rendered images of our generated SVG and this professional design collection, using features extracted by the CLIP image encoder [radford2021learning].
Path-level Regularity. We assess path quality using a transformer-based Auto-Encoder trained on the FIGR-8-SVG [clouatre2019figr] dataset, following the DeepSVG [carlier2020deepsvg] architecture. Our model, trained with reconstruction tasks, encodes each drawing command into a latent vector. We represent each path as the mean of its commands’ latent embeddings, and calculate the FID between these path representations and the ground truth paths from FIGR-8-SVG. This metric quantifies how closely our generated paths align with professional SVG design patterns.
Text-level Alignment. We assess semantic alignment by computing the CLIP score [radford2021learning] between the text prompt and the rendered SVG image.
@@ 4.1 Comparison with Existing Methods
| CLIPDraw [frans2021clipdraw] | |||||||||
| DiffSketcher [xing2023diffsketcher] | |||||||||
| VectorFusion [jain2022vectorfusion] | |||||||||
| SVGDreamer [xing2024svgdreamer] | |||||||||
| T2V-NPR [zhang2024text] | |||||||||
@@
Qualitative Comparison. In Figure 3, we present a side-by-side comparison between our Chat2SVG and the baselines. Methods that optimize open-ended strokes based on CLIP or Diffusion models (i.e., CLIPDraw [frans2021clipdraw] and DiffSketcher [xing2023diffsketcher]) produce results with messy visual appearances and poor text alignment. These methods sometimes fail to synthesize complete objects, such as the flamingo (first row) and the plane (third row) in Figure 3. In contrast, our method generates SVG results with clear layouts and greater fidelity to the prompts.
Methods that optimize closed shapes via score distillation of diffusion models (i.e., VectorFusion [jain2022vectorfusion], SVGDreamer [xing2024svgdreamer], and T2V-NPR [zhang2024text]) can better align with the input text, but still produce fragmented paths with limited semantic meaning. For instance, in the plane example (third row of Figure 3), VectorFusion and SVGDreamer create planes and clouds composed of multiple jagged, irregular, and fragmented shapes, which only appear meaningful when viewed as a whole. This hinders convenient editing by graphic designers. T2V-NPR [zhang2024text] addresses the issue of jagged paths using a path VAE (although it can be overly smooth at times) and reduces fragmented shapes by merging shapes with similar colors. However, its merging operation ignores the semantic meaning of shapes, resulting in the body of the plane being merged with the surrounding clouds. Additionally, the semantic meaning of the path is sometimes unclear, as exemplified by the cloud in the bottom-right corner, whose contour is actually approximated by the blue background instead of a true cloud shape. In contrast, our Chat2SVG adopts a unique approach by using LLMs to generate paths representing semantic components and performing dual-stage optimization based on image diffusion to enhance path expressiveness, thus ensuring both path regularity and visual fidelity of generated SVG.
@@ 5 Conclusion
Despite our method’s superior results, there are several aspects that could be further improved: (1) LLMs may produce imprecise visual layouts, such as the burger in Figure 7. This can be alleviated by collecting an SVG dataset and extracting the corresponding layouts to fine-tune LLMs. (2) Our approach uses SAM to identify new parts. However, SAM may overlook important regions, like the pink frosting on the donut in Figure 7. Incorporating semantic-aware SAM [li2023semantic] could potentially produce more accurate results. (3) The pretrained SVG VAE from [zhang2024text] has a fixed number of commands. This makes it hard to represent intricate shapes. For example, the lion’s mane in Figure 7 lacks sufficient detail. Training a new SVG VAE with varying-length commands would help solve this problem. Our work can be extended to generate other types of vectorized data, such as icons and fonts. Additionally, through consistent editing of SVG across multiple frames, coupled with shape interpolation, our Chat2SVG has the potential to facilitate keyframe animation generation.
Acknowledgement
The work described in this paper was fully supported by a GRF grant from the Research Grants Council (RGC) of the Hong Kong Special Administrative Region, China [Project No. CityU 11216122].
@@ Appendix A Overview
More generated and edited SVG examples (Section E) are available on the project page.
@@ Appendix E More Results
To provide a more user-friendly visualization, we have put all generated SVG, including the text-guided generation and editing results, on the project website. Please visit the https://chat2svg.github.io/ to view more outputs.
[…]